Find the value of b for which the equation

(ye^2xy+x)dx +bxe^ 2xy dy= 0



1
Expert's answer
2021-06-09T08:23:56-0400

Let us find the value of "b" for which the equation "(ye^{2xy}+x)dx +bxe^{2xy} dy= 0" is an exact differential equation. Since "\frac{\partial(ye^{2xy}+x)}{\partial y}=e^{2xy}+2xye^{2xy}" and "\frac{\partial(bxe^{2xy})}{\partial x}=be^{2xy}+2bxye^{2xy}", we conclude that "\frac{\partial(ye^{2xy}+x)}{\partial y}=\frac{\partial(bxe^{2xy})}{\partial x}" implies "e^{2xy}+2xye^{2xy}=be^{2xy}+2bxye^{2xy}", and hence "b=1."

Answer: "b=1."